OpenStudy (anonymous):

if the function f,g and h are defined by f(x)=2x , g(x)=cosx , and h(x)=x/3 , then f(g(h(22.3/2)))=?

OpenStudy (shamim):

f(g(h(22.3/2)))=f(g(22.3*3/2))=f(cos(22.3*3/2))= 2cos(22.3*3/2)

OpenStudy (shamim):

did u get it